HUBBLE SPACE TELESCOPE DAILY REPORT #2102 PERIOD COVERED: 0000Z (UTC) 04/14/98 - 0000Z (UTC) 04/15/98 Daily Status Report as of 105/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD/NIC/3/WF/PC-2 7355 (Continuous Ultraviolet Monitoring of NGC 3516) The Space Telescope Imaging Spectrograph (CCD), the Near Infrared Camera (Camera 3), and the WF/PC-2 were used to make continuous ultraviolet observations of galaxy NGC-3516. Recent intensive IUE monitoring of NGC 4151 and other AGN reveal rapid aperiodic variations down to the shortest time scales yet probed {~80 min}, strongly correlated throughout the UV and X-ray with no detectable interband lag. A wavelength calibration image was also taken. As described in 2.1, two of the re-acquisitions for this proposal experienced at least four losses of lock and ultimately failed, affecting four NICMOS observations, eight STIS observations, and four WF/PC-2 observations. Otherwise, the observations were completed as planned, and no anomalies were reported. 1.2 Completed WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 5 Successful : 3 As documented in HSTAR 6436, the re-acquisitions at 104/2225Z and at 104/2355Z failed after at least four losses of lock. The proposal outlined in 1.1 was affected. 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Operations Notes: The SSR correctable EDAC counts were cleared five times, per ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH